OpenStudy (anonymous):

Ingrid is making a quilt using squares that measure 5 in. on a side. What is the length of a diagonal of one of the quilt squares? Round to the nearest tenth.

OpenStudy (anonymous):

well from what i interpret is that its a square and you want to find the diagonal, which would look a bit like this;  so i think too solve it we use the Pythagorean theorem, which is a^2 + b^2 = c, where c is the hypotenuse which is the diagonal, and a and b are the sides. 5^2 + 5^2 = c^2 25 + 25 = c^2 50 = c^2 c^2 = 50 c = *sqrt(50 c = 5*sqrt(2 but since it says to round to the nearest tenth, using the calculator we find out the decimal number for it which is 7.0710; we round it and it ends up as 7.1